OpenStudy (anonymous):

find the volume of 294 cm^2 in m^2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no :c

OpenStudy (skullpatrol):

Do you know what the prefix "centi" means in centimeter?

OpenStudy (anonymous):

No i don't

OpenStudy (campbell_st):

well think of it this way 1 metre = 100 centimetres so a square 1 metre by 1 metre has an area 100cm by 100cm or 100 x 100 = then if you have a cube 1 metre high the volume is 100 x 100 x 100 = cm^3

OpenStudy (anonymous):

okay so because its cm^2 it would be 100 x 100 ?

OpenStudy (campbell_st):

well I'm not sure you have asked about volume but the units given are area...?

OpenStudy (anonymous):

Honestly am not very sure what this problem is asking to do

OpenStudy (anonymous):

one thing for sure its asking for the volume

OpenStudy (campbell_st):

is its area in cm to metres is \[A = \frac{294}{100 \times 100} m^2\]

OpenStudy (skullpatrol):

@campbell_st why did you talk about volume?

OpenStudy (campbell_st):

because the question talks about volume

OpenStudy (skullpatrol):

But the units are for area.

OpenStudy (campbell_st):

that was my question, the question says find the volume of 294 cm^2 in m^2 the question asks for volume but gives the units as area.

OpenStudy (skullpatrol):

So there must be a mistake in the question.

OpenStudy (campbell_st):

yep... thats what I've asked.

OpenStudy (anonymous):

well to give you more information the title of the unit it "Changing Units of Measure"

OpenStudy (campbell_st):

I understand the question @Jen_jen95 can you check is Volume and the number is 294 cm^2 and not 294 cm^3

OpenStudy (skullpatrol):

Since ^2 is mentioned twice, I believe they meant Area, not volume.

OpenStudy (skullpatrol):

$$ centi = 10^{-2} = \frac{1}{100}$$ @Jen_jen95 have you seen this conversion factor before?

OpenStudy (anonymous):

only for cube

OpenStudy (skullpatrol):

What do you mean?

OpenStudy (anonymous):

actually i have it ! c: thank you for your help c: <3

OpenStudy (skullpatrol):

Thanks for trying c: